### v2

# Automotive Speed Estimation: Sensor Types and Error Characteristics from OBD-II to ADAS

Abstract
Modern on-road navigation systems heavily depend on integrating speed measurements with inertial navigation systems (INS) and global navigation satellite systems (GNSS). Telemetry-based applications typically source speed data from the On-Board Diagnostic II (OBD-II) system. However, the method of deriving speed, as well as the types of sensors used to measure wheel speed, differs across vehicles. These differences result in varying error characteristics that must be accounted for in navigation and autonomy applications. This paper addresses this gap by examining the diverse speed-sensing technologies employed in standard automotive systems and alternative techniques used in advanced systems designed for higher levels of autonomy, such as Advanced Driver Assistance Systems (ADAS), Autonomous Driving (AD), or surveying applications. We propose a method to identify the type of speed sensor in a vehicle and present strategies for accurately modeling its error characteristics. To validate our approach, we collected and analyzed data from three long real road trajectories conducted in urban environments in Toronto and Kingston, Ontario, Canada. The results underscore the critical role of integrating multiple sensor modalities to achieve more accurate speed estimation, thus improving automotive navigation state estimation, particularly in GNSS-denied environments.
Index Terms:
Navigation Technologies, Instrumentation, On-board Diagnostics, Speed Estimation, Error Modeling.
I Introduction
Accurate speed estimation, particularly forward velocity, is crucial for automotive applications such as fleet management, navigation, Advanced Driver Assistance Systems (ADAS), and Autonomous Driving (AD) [1]. Vehicle tracking depends on real-time vehicular data for route optimization, while ADAS and AD require precision for functions like lane-keeping and collision avoidance. Mercedes-Benz’s DRIVE PILOT [2] emphasizes redundancy and safety, using moisture sensors in the wheel wells to monitor road wetness and backup systems to ensure reliable operation in challenging conditions.
Many complementary systems, such as navigation and driver assistance solutions added post-manufacture, access essential vehicular data through the On-Board Diagnostic II (OBD-II) port [3]. As illustrated in Fig. 1, the OBD-II interface connects to the vehicle’s primary network, typically the CAN bus, or other protocols depending on the make and model. It allows communication with subsystems such as the Electronic Control Units (ECUs), Anti-lock Braking System (ABS), Instrument Cluster, and Adaptive Cruise Control (ACC) modules, to name a few. This standardized interface provides critical parameters, including engine RPM and wheel speed, facilitating seamless integration with internal and external systems. However, the sensors and techniques used to derive speed measurements can differ significantly between land vehicles. These differences can introduce variations in error characteristics, which can profoundly impact the reliability and accuracy of downstream systems relying on this data.
Figure 1: Interconnections of Automotive Sensors and Systems via CAN.
Wheel speed, in particular, plays a critical role in enhancing Inertial Navigation Systems (INS) and enabling the Reduced Inertial Sensor System (RISS), a simplified INS variant that predominantly depends on external speed inputs for operation [4]. The quality of wheel speed data is essential in GNSS-denied environments or, in the case of modern systems, signals of opportunity (SOPs)-denied conditions, where errors in proprioceptive measurements can propagate and degrade dead-reckoning performance. Furthermore, a limited understanding of sensor types, error characteristics, and data formats can compromise the accuracy and reliability of complementary systems. The key contributions of this paper are:
Figure 2: Methods for estimating the speed of a land vehicle without relying on GNSS or IMU integration.
  1. 1)
    Comprehensive overview of automotive speed sensing: We analyze methods for estimating vehicle speed, highlighting classical and modern sensor technologies.
  2. 2)
    Comparison of speed derivation processes: We investigate the speed derivation methods employed at the Original Equipment Manufacturer (OEM) level for different vehicle types. We also emphasize the distinctions between OBD-II and CAN bus protocols’ data format.
  3. 3)
    Real-road data collection and analysis: We utilize road test data to identify speed sensor types from OBD-II-reported messages, compare the accuracy of OBD-II-derived speed estimates with perception-based methods, and conduct an error variance analysis as a step toward achieving effective sensor fusion.
The remainder of this paper is organized as follows: Section II reviews related work on ground vehicle speed sensing and its applications in navigation. Then, Section III explores OBD-II communication, data formatting, and speed derivation for various vehicle configurations, along with an overview of wheel speed sensors (WSS). Following this, Section IV presents experimental results, including sensor identification, error variance analysis, and the integration of stereo cameras for complementary speed estimation. Finally, Section V concludes the paper and outlines future research directions.
II Related Work
Several publicly available automotive datasets offer valuable insights for research and development; however, only a few include data collected directly through a vehicle’s proprietary diagnostic interface. Among datasets that provide speed data, high-resolution Distance Measuring Instruments (DMIs), such as those featured in the DARPA dataset [5] and the Ford Campus dataset [6], are commonly used. While these systems deliver exceptional accuracy, their high cost and complexity often exceed the requirements of low-cost automotive navigation applications.
In addition to built-in vehicular speed sensing technologies and DMIs, alternative systems have been explored for specialized ground vehicle testing. Magnetic Rotary Encoders (MREs) [7] and Non-contact Optical Sensors [8, 9] are examples of technologies used in extreme conditions, such as snow, ice, or wet surfaces, or in harsh environments characterized by dust, moisture, extreme temperatures, and mechanical shocks. Despite their robustness and precision, these niche applications and high costs limit their practical use in everyday automotive systems.
Datasets like nuScenes [10], A2D2 [11], and comma2k19 [3] aim to make built-in vehicular data more accessible. These datasets rely on the CAN bus network, which poses implementation challenges due to the lack of standardization across manufacturers. Each manufacturer defines unique CAN message structures and IDs, typically documented in DBC files. Consequently, interpreting data such as vehicle wheel speed from the CAN bus often requires extensive knowledge of the vehicle’s architecture or reverse engineering. In contrast, the OBD-II protocol provides a standardized and interoperable approach for extracting diagnostic data, including vehicle speed, across all OBD-II-compliant vehicles. Table I summarizes the key differences between the OBD-II protocol and CAN bus. Although the OBD-II protocol is straightforward, the diverse methods used for deriving vehicle speed from built-in sensors remain underexplored in the literature, creating significant challenges for navigation system integrators aiming to develop universally compatible, robust, and low-cost solutions.
TABLE I: Comparison between OBD-II protocol and CAN Bus
Aspect OBD-II Standard CAN Bus
Purpose Diagnostics and emissions monitoring Real-time communication between ECUs
Standardization Standardized PIDs (e.g., speed: 0x0D) Manufacturer-specific (e.g., CAN ID varies)
Data Access Limited to a subset of vehicle parameters Full access to all CAN messages
Communication Model Request-Response Broadcast (continuous streaming)
Speed of Access Slower, query-based Faster, real-time
Physical Access OBD-II port OBD-II port (or direct access to CAN wires)
In addition to the often omitted derivation process, the navigation and instrumentation literature offers limited clarity on the specific sensors automotive suppliers use for speed estimation. Researchers often overlook the type of speed sensors employed in their test vehicles, commonly employ the general term odometer which covers a wide variety of instruments, and frequently rely on traditional Inertial Measurement Unit (IMU) and GNSS error modeling techniques for sensor fusion [4, 12, 13, 14]. While these approaches have been widely used, they may not guarantee optimal estimation of vehicle speed. Therefore, identifying and characterizing vehicular speed sensors is critical for designing robust and effective navigation systems. Fig. 2 categorizes vehicle speed-sensing technologies into two primary groups: built-in diagnostic data and alternative sensors and systems. The first group includes OBD-II-derived speed measurements and CAN bus-derived data from Vehicle Speed Sensors (VSS) and WSS, both of which are widely integrated into vehicles’ onboard systems. The second group consists of external sensors, such as non-contact optical sensors, DMIs, and perception-based systems like cameras, LiDARs, and RADARs. Cameras and RADARs, in particular, are increasingly utilized not only for ADAS but also for forward velocity estimation. Recent studies have demonstrated the effectiveness of cameras for ego-motion tracking [15] and RADARs for ego-velocity estimation [16]. These perception-based approaches offer practical and cost-effective alternatives, particularly in scenarios where traditional automotive speed sensors are limited or unavailable.
III OBD-II Vehicle Speed Derivation
Modeling errors in consumer-level OBD-II-derived speed measurements presents a significant challenge due to the restricted data available through the vehicle’s regular OBD-II protocol. Generally, the consumer-level OBD-II provides data on the overall vehicle speed rather than the speed of each individual wheel. Additionally, the speed data is often in integer format, which lacks the resolution of decimal points found in double or float types commonly available in CAN bus systems, leading to less precise measurements. When querying the OBD-II system for vehicle speed, a command is sent, and the system responds with encoded data. Below is an example illustrating this process:
1# Command Sent
2> 010D
3# Response Received
441 0D 3C
5Interpretation:
6# 41: Response for mode 01
7# 0D: Echo of the PID for vehicle speed
8# 3C: Vehicle speed in hexadecimal (60 in decimal, km/h)
Listing 1: OBD-II Command and Response Example.
The command 010D represents a request for the current vehicle speed (PID 0x0D). The response 41 0D 3C indicates the vehicle speed as 0x3C, which is 60 $km/h$ when converted to decimal [^1] Automakers may configure ECUs to report diagnostic data in region-specific units. For example, vehicle speed may be in miles per hour ($mph$) for the USA or in kilometers per hour ($km/h$) for Canada and Europe.. In other words, the reported speed through OBD-II interfacing is:
$V_{\text{v}}^{\text{OBD}}=\left\lfloor\hat{V}\right\rceil.$ (1)
[…]
III-A Transmission-based Vehicle Speed Determination
The estimation of vehicle speed ($V_{\text{v}}$) from transmission data relies on the relationship between engine speed ($N_{e}$) in RPM and drivetrain parameters, especially gear ratio. The gear ratio governs the conversion of engine speed (input) to wheel speed (output). For example, in Fig. 3, a driving gear with 8 teeth and a driven gear with 16 teeth creates a 2:1 ratio, where the output rotates at half the input speed. This adjustment is crucial for controlling speed and torque. The gearbox alters engine speed based on the engaged gear ratio ($i_{x}$), which changes with each gear shift, while the differential uses its gear ratio ($i_{0}$) to balance wheel speeds, critical for smooth turning.
Figure 3: Illustration of gear ratio.
In non-ABS vehicles, where speed is derived from the engine’s VSS (see Fig. 4), the calculation for speed in $m/s$ is given by the following formula:
$V_{\text{v}}=\left\lfloor\left({\frac{N_{e}\cdot\pi\cdot r_{w}}{30\cdot i_{x}\cdot i_{0}}}\right)\right\rceil,$ (2)
$N_{e}$ represents the engine speed, typically measured by a crankshaft position (CKP) sensor or an engine speed sensor. The wheel radius ($r_{w}$), measured in meters ($m$), converts the rotational speed of the wheels into linear speed. This calculation assumes a slip-free mechanical connection between the engine and wheels, ensuring efficient power transfer and accurate speed estimation.
Figure 4: Vehicle Speed Sensor (VSS).
III-B ABS-based Vehicle Speed Determination
The ABS WSS, depicted in Fig. 5, produces output signals as the wheel turns, i.e. wheel ticks. These signals are then transmitted to the vehicle’s ECU or a specific ABS control module. The main role of the ECU or ABS module is to interpret these signals to determine each wheel’s speed.
Figure 5: Wheel Speed Sensor (WSS).
The compute module calculates the speed of each individual wheel using the formula:
$V_{\text{w}}=\frac{2\pi\cdot r_{w}\cdot N_{p}}{N_{t}\cdot\Delta t},$ (3)
where $V_{\text{w}}$ represents the speed of a single wheel in $m/s$, $r_{\text{w}}$ denotes the radius of the wheel in $m$, $N_{p}$ is the count of pulses detected within the time interval $\Delta t$, and $N_{t}$ is the total number of teeth on the reluctor ring. To determine the vehicle’s overall speed, the ECU averages the rotational speeds of all the individual wheels. This average is then used for various control and monitoring functions within the vehicle’s ABS and electronic stability control systems (i.e., ESC).
Compared to the transmission-based system, ABS-based speed estimation offers greater robustness and accuracy while involving fewer parameters. Thus, mathematically, the actual ABS wheel speed reported through the regular consumer-level OBD-II protocol is computed as follows:
$V_{\text{v}}=\left\lfloor\left(\frac{V_{\text{w}_{1}}+V_{\text{w}_{2}}+V_{\text{w}_{3}}+V_{\text{w}_{4}}}{4}\right)\right\rceil,$ (4)
where $V_{\text{v}}$ is the speed of the vehicle, $V_{\text{w}_{i}}$ is the velocity of the $i^{th}$ wheel in $m/s$, assuming each wheel has an installed WSS. To overcome these challenges without relying on automotive OEM collaboration or specialized tools (i.e., reverse engineering the CAN bus) to access detailed information about wheel sensors in test vehicles, navigation, and mapping entities frequently employ DMIs.
III-C Wheel Speed Sensor Types
When dealing with WSS, limited information can be found in the navigation and instrumentation literature related to the type of WSS automotive suppliers use and their corresponding accuracy. Within the automotive industry, two primary types of wheel speed sensors are currently utilized, each possessing unique advantages and disadvantages. These are referred to as $passive$ and $active$ wheel speed sensors, respectively. The operational mechanisms of these sensors exceed the scope of this paper, however, their main differences are shown in Table II.
TABLE II: Automotive WSS: A comparison.
WSS Types
Passive Active
Advantages - Lower cost - Better accuracy at
low speeds
- Easy to maintain - Suitable for high-
performance applications
- No power supply required - Advanced diagnostic
capabilities
Drawbacks - Poor accuracy at - Higher cost
low speeds - More complex design
- Limited diagnostic - More susceptible to
capabilities electromagnetic interference
While Passive WSS (PWSS) are simple in terms of design and operation, passive sensors may be less accurate, especially at lower speeds. In fact, the ECU typically disregards speed signals below a certain speed threshold ($\leq$ 3 $km/h$) to prevent false triggering from signal noise. In contrast, active WSS can detect speeds as low as $0.1$ $km/h$ [17] and deliver consistent error characteristics across the vehicle speed spectrum under ideal conditions due to their advanced design and operation.
IV Experimental Work and Results
This section describes the experimental setup, details of the road tests, and results from three real-road trajectories performed in Kingston and Toronto, Ontario, Canada.
IV-A Experimental Setup
The experiments were conducted using an ABS-equipped Toyota Sienna minivan, with the logging setup briefly depicted in Fig. 6. The dataset collected during these experiments has been made publicly available [^2] The dataset can be accessed at: https://github.com/hanymragab/navinst19.
Figure 6: Data logging setup.
[…]
IV-B Results and Discussion
IV-B1 Identification and Error Modeling of Automotive Speed Sensors
In integrated navigation systems, understanding how the integrity of sensed measurements impacts overall navigation performance is essential. Consequently, performing error variance analysis on logged vehicle speed measurements is a critical step in evaluating sensor accuracy. To achieve reliable error variance estimates, real-world data must be collected from multiple road test trajectories conducted with the same vehicle. The data from each trajectory can then be post-processed to estimate the average error variance as a function of a quantity known to influence measurement accuracy.
This average error variance can be modeled using a suitable mathematical function, such as a polynomial or exponential fit, to characterize its behavior with respect to the influencing quantity. Such modeling provides valuable insights into the error characteristics, enabling more informed decisions for the design and optimization of sensor fusion algorithms and related systems.
Figure 7: Error variance analysis on passive wheel speed sensor data gathered throughout three different trajectories.
Fig. 7 shows the results for the data collected in the cities of Kingston and Toronto that were used to fit the error variance. Upon initial data inspection, the quantity of interest that was found to impact measurement accuracy was the speed of travel. This is a clear indication that the sensor used is $passive$ in its nature. Moreover, the reason why the 0-4 $km/h$ speed interval was omitted in the variance analysis is simply due to the fact that PWSS may not report speeds lower than around 3 $km/h$ as shown in Fig. 8.
Figure 8: OBD-II-reported vehicle speed (green) as compared to the reference high-end solution (red).
The vehicular forward speed measurements were directly compared against the resultant velocity computed from the high-end tactical-grade integrated navigation solution. After further data analysis, a 2nd degree exponential function was selected in the data fitting process as it provides a flexible and efficient way to model the data accounting for non-linearities. The 2nd degree exponential function takes the form of:
$\sigma^{2}_{PWSS}(v_{f})=ae^{bv_{f}}+ce^{dv_{f}},$ (5)
where the coefficients $a$, $b$, $c$, and $d$, are determined using an optimization algorithm. In this case, the obtained function is $\sigma^{2}_{PWSS}(v_{f})=0.3095e^{-0.1241v_{f}}+0.1477e^{-0.0009149v_{f}}$ and the coefficients are determined with 95% confidence bounds. This means that there is a 95% probability that the actual values of the coefficients fall within the bounds of the confidence intervals which corresponds to ±$\pm 3\sigma$. The obtained function provides a good representation of the data and can be used for further analysis or predictions.
Figure 9: PDF in terms of speed intervals for three real road trajectories, two of which are in Toronto, Ontario, Canada.
Driving patterns vary significantly depending on the environment, with vehicles typically traveling slower in dense urban areas and downtown cores compared to highways and other parts of the road network. Fig. 9 illustrates this trend, showing that the probability of a car traveling at speeds between 0–4 $km/h$ is 35–45% for trajectories T19.1 and K19.1, which traverse urban and downtown cores. In contrast, for T19.2, which includes a significant portion of highways, this probability drops below 30%. In dense, GNSS-denied urban areas, evaluating perception systems for speed estimation becomes particularly relevant.
IV-B2 OBD-II-derived speed vs Stereo VO-derived speed
To evaluate the performance of speed estimates across different intervals, we compare OBD-II-derived vehicle speeds with results from a stereo visual odometry (Stereo VO) routine and an enhanced Stereo VO algorithm incorporating semantic segmentation-based outlier rejection (Stereo VO-SS-OR), as introduced in our earlier work [18]. Fig. 10 reveals significant variations in the accuracy of speed measurements. In the lower speed range (0–4 $km/h$), both Stereo VO algorithms outperform the OBD-derived PWSS speed measurements, with Stereo VO-SS-OR achieving the lowest root mean square error (RMSE), unlike at higher speed intervals.
Figure 10: Speed RMSE comparison between PWSS and Stereo VO with and without SS-OR.
Fig. 11 presents a comparison of speed estimates from Stereo VO, with and without the SS-OR enhancement, alongside OBD-II-derived speeds and the NovAtel reference. It highlights the variability in Stereo VO performance across different scenarios and showcases the capability of the SS-OR-enhanced version to improve vehicular speed estimation, particularly when PWSS fails to provide data. The integration of SS-OR significantly enhances ego-motion estimation accuracy, resulting in more consistent and precise speed measurements, especially at lower speeds of travel in highly dynamic environments.
Figure 11: Performance at low travel speeds.
On the other hand, Fig. 12 illustrates the challenges faced in low-traffic environments dominated by static objects, which significantly affect the approach’s effectiveness. In such scenarios, Stereo VO generates noisier and less accurate speed estimates compared to OBD-II-derived speeds. The performance of Stereo VO-SS-OR deteriorates further at higher travel speeds, primarily due to baseline discrepancies
Figure 12: Degeneration of SS-OR: Parked or stationary vehicles misclassified as dynamic, with suboptimal travel speed (vs framerate) and illumination.
between frames combined with framerate limitations. Additionally, the abundance of parked cars (as highlighted in the figure), often being misclassified as dynamic within the SS-OR routine, disrupts temporal feature tracking, and impairs accurate ego-motion estimation. Given these differences, implementing a fusion scheme is essential to selectively weight measurements based on WSS speed data, enabling a more accurate estimation of forward speed. Additionally, this highlights the critical need for integrating object tracking or supplementary sensors, such as RADAR sensors to improve stationary and dynamic object classification, enhancing the overall performance of the SS-OR framework.
V Conclusion
In conclusion, this paper sheds light on a critical yet underexplored aspect of navigation and instrumentation: the diverse sensors and methods of deriving automotive speed measurements from the OEM perspective. By identifying and analyzing these approaches, we demonstrated the strengths and limitations of using OBD-II-derived speed data for land vehicle navigation. Through experiments conducted on three extended trajectories, we identified the wheel speed sensor type employed and showcased the advantages of using alternative ADAS sensors, such as front-facing stereo cameras, for speed estimation.
While each sensor modality has its own strengths and limitations, our findings emphasize the necessity of a fusion engine to combine data sources effectively, ensuring a robust integration with INS or RISS systems. To facilitate universal plug-and-play functionality, it would be beneficial to gather OBD-II-derived speed data from a wide range of vehicles, including those equipped with active wheel speed sensors. This approach would enable the development of more accurate error profiles and an adaptive system that enhances the fusion process with low-cost exteroceptive sensors.
Looking ahead, for more sophisticated and car-specific systems, reverse engineering the CAN bus to access individual wheel speed data directly presents a promising avenue. This approach has the potential to significantly enhance the accuracy of speed estimation, ultimately leading to substantial improvements in overall navigation system performance.